# Characteristic Learning for Provable One Step Generation

### v3

@@ 4.2. Experiment results and discussions
In this section, we validate the generation quality and sampling efficiency of the characteristic generator using both synthetic and real data through numerical experiments. Föllmer flow is token as the underlying ODE, and all results can be generalized to arbitrary probability flow ODE without loss of generality. We use the Fréchet inception distance (FID) to measure sampling quality on image data. Lower FID means better performance.
@@ 4.2.2. Real dataset
We compare the images generated by the numerical sampler and characteristic generator with different numbers of function evaluations (NFE) in Figure 4. The numerical sampler fails to accurately generate images at small NFE values such as 1 and 2. In fact, with 1 NFE, the solution is actually close to the mean of the target distribution. It is necessary to have five or more NFE for the numerical ODE solvers to work properly. In contrast, the characteristic generator is capable of generating high-quality images even with only 1 NFE.
@@ 6. Conclusions
On the theoretical front, we intend to exploit the regularity of the velocity field in our error analysis for velocity matching. This will allow us to achieve a faster convergence rate. Additionally, we plan to analyze higher order and more stable numerical schemes, such as the high-order exponential integrator, in order to provide a comprehensive understanding of their effectiveness. Furthermore, we aim to establish a theoretical foundation for the role of semi-group penalties in the characteristic fitting. This will contribute to a deeper understanding of their impact and significance in our framework.
Appendix A Supplemental Definitions and Lemmas
@@
We first introduce sub-Gaussian random variable [Vershynin, 2018, Wainwright, 2019].
Definition A.1 (Sub-Gaussian).
@@
The following results (Lemmas A.2 to A.4) shows some important properties of sub-Gaussian variables, whose proofs can be found in Wainwright [2019, Section 2.1].
Lemma A.2 (Chernoff bound).
@@
Lemmas A.5 and A.6 show bounds of the tail probability and the expectation of the maximum of $N$ sub-Gaussian variables.
Lemma A.5.
@@
where the last inequality holds from Lemma A.2. This completes the proof. ∎
Proof of Lemma A.6.
@@
where the last inequality holds from Lemma A.3 for each $\lambda\in[0,1)$. Letting $\lambda=1/2$ yields the desired inequality. ∎
Lemma A.7 (Fourth moment of standard Gaussian).
@@
We next introduce the notation of covering number and Vapnik-Chervonenkis dimension (VC-dimension), both of which measure the complexity of a function class [Mohri et al., 2018, Vaart and Wellner, 2023]. They are used in the error analysis for velocity matching (Section D) and characteristic fitting (Section F).
Definition A.8 (Covering number).
@@
Definition A.9 (VC-dimension).
Let ${\mathscr{F}}$ be a class of functions from ${\mathbb{R}}^{d}$ to {±$\{\pm 1\}$. For any non-negative integer $m$, we define the growth function of ${\mathscr{F}}$ as
@@
A set $\{X^{(i)}\}_{i=1}^{m}$ is said to be shattered by ${\mathscr{F}}$ when |$|\{(f(X^{(1)}),\ldots,f(X^{(m)})):f\in{\mathscr{F}}\}|=2^{m}$. The Vapnik-Chervonenkis dimension of ${\mathscr{F}}$, denoted $\operatorname{VCdim}({\mathscr{F}})$, is the size of the largest set that can be shattered by ${\mathscr{F}}$, that is, $\operatorname{VCdim}({\mathscr{F}})=\max\{m:\Pi_{{\mathscr{F}}}(m)=2^{m}\}$. For a class ${\mathscr{F}}$ of real-valued functions, we define $\operatorname{VCdim}({\mathscr{F}})=\operatorname{VCdim}(\operatorname{sign}({\mathscr{F}}))$.
Lemma A.10.
Let ${\mathscr{F}}$ be a class of functions mapping from ${\mathbb{R}}^{d}$ to ${\mathbb{R}}$, and let ${\mathscr{H}}$ be a function class defined as =${\mathscr{H}}=\{(x,f)\mapsto h(f,x)\in{\mathbb{R}}:f\in{\mathscr{F}}\}$. Suppose ${\mathcal{D}}=\{X^{(i)}\}_{i=1}^{n}$ is a set of samples in ${\mathbb{R}}^{d}$. If there exists a constant $L>0$ such that for each $f,f^{\prime}\in{\mathscr{F}}$,
@@
Lemma A.11 (Anthony et al. [1999, Theorem 12.2]).
@@
Lemma A.12 (Bartlett et al. [2019, Theorem 7]).
@@
Appendix B Proof of Results in Section 2
The proofs of Propositions 2.1 and 2.3 are shown in this section.
The proof of Proposition 2.1 follows from the proof of Albergo et al. [2023b, Theorem 2] and Albergo et al. [2023a, Theorem 2.6].
@@ Appendix B Proof of Results in Section 2
(B.1) $\frac{\partial}{\partial t}\Big{(}\int_{{\mathbb{R}}^{d}}\phi(x)\rho_{t}(x)\,\mathrm{d}x\Big{)}=\int_{{\mathbb{R}}^{d}}\phi(x)\partial_{t}\rho_{t}(x)\,\mathrm{d}x.$
@@
(B.2) $=\int_{{\mathbb{R}}^{d}}b_{t}^{*}(x)\cdot\nabla\phi(x)\rho_{t}(x)\,\mathrm{d}x=-\int_{{\mathbb{R}}^{d}}\phi(x)\nabla\cdot(b_{t}^{*}(x)\rho_{t}(x))\,\mathrm{d}x,$
where the first equality follows from the chain rule, the forth equality holds from the definition of $b_{t}^{*}$ (2.3), and the last equality is due to integration by parts and the divergence theorem [Evans, 2010, Theorem 1 in Section C.2]. Combining (B.1) and (B.2) gives the desired transport equation. ∎
Proof of Proposition 2.3.
See Lemma D.3 and its proof for details. ∎
Appendix C Properties of the Probability Flow ODE
In this section, we present some auxiliary properties of the probability flow ODE in Section C.1. Proofs of the propositions in Section 3.2 are given in Section C.2.
C.1. Auxiliary Properties
Lemmas C.1 to C.4 are established under Assumption 1.
Lemma C.1.
@@ C.1. Auxiliary Properties
where the we used the definition of conditional density and Lemma C.1. This completes the proof. ∎
Lemma C.3.
@@
$b^{*}(t,x)=\frac{\dot{\alpha}_{t}}{\alpha_{t}}x+\beta_{t}\Big{(}\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\Big{)}{\mathbb{E}}[X_{1}|X_{t}=x],$
@@
(C.1) $\nabla_{x}b^{*}(t,x)=\frac{\dot{\alpha}_{t}}{\alpha_{t}}I_{d}+\beta_{t}\Big{(}\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\Big{)}\nabla_{x}{\mathbb{E}}[X_{1}|X_{t}=x].$
@@
(C.2) $=-\nabla_{x}\log\rho_{t}(x)\int_{{\mathbb{R}}^{d}}x_{1}^{T}\rho_{1|t}(x_{1}|x)\,\mathrm{d}x_{1}+\int_{{\mathbb{R}}^{d}}\nabla_{x}\log\rho_{t|1}(x|x_{1})x_{1}^{T}\rho_{1|t}(x_{1}|x)\,\mathrm{d}x_{1},$
@@
For the first term in the right-hand side of (C.2), applying Lemma C.2 implies
@@
(C.3) $=\Big{(}\frac{1}{\alpha_{t}^{2}}x-\frac{\beta_{t}}{\alpha_{t}^{2}}{\mathbb{E}}[X_{1}|X_{t}=x]\Big{)}{\mathbb{E}}[X_{1}|X_{t}=x]^{T}.$
For the second term, it follows from Lemma C.1 that
@@
(C.4) $=-\frac{1}{\alpha_{t}^{2}}x{\mathbb{E}}[X_{1}|X_{t}=x]^{T}+\frac{\beta_{t}}{\alpha_{t}^{2}}{\mathbb{E}}[X_{1}X_{1}^{T}|X_{t}=x].$
Plugging (C.3) and (C.4) into (C.2) yields
(C.5) $\nabla{\mathbb{E}}[X_{1}|X_{t}=x]=\frac{\beta_{t}}{\alpha_{t}^{2}}\operatorname{Cov}(X_{1}|X_{t}=x).$
Substituting (C.5) into (C.1) completes the proof. ∎
@@
Lemmas C.5 to C.8 are established under Assumptions 1 and 2.
Lemma C.5.
@@
$(X_{1}|X_{t}=x)\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}U+\sqrt{\frac{\sigma^{2}\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}}\epsilon+\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}x.$
Proof of Lemma C.5.
@@
(C.6) $\rho_{t|1}(x|x_{1})=\frac{1}{(2\pi)^{d/2}\alpha_{t}}\exp\Big{(}-\frac{1}{\alpha_{t}^{2}}\|\beta_{t}x_{1}-x\|_{2}^{2}\Big{)}.$
@@
(C.7) $\rho_{1}(x_{1})=\frac{1}{(2\pi)^{d/2}\sigma}\int_{{\mathbb{R}}^{d}}\exp\Big{(}-\frac{1}{\sigma^{2}}\|x_{1}-u\|_{2}^{2}\Big{)}\,\mathrm{d}\nu(u).$
Combining (C.6) and (C.7) implies

### v5

@@ 4.2. Experiment results and discussions
In this section, we validate the generation quality and sampling efficiency of the characteristic generator using both synthetic and real data through numerical experiments. Föllmer flow is taken as the underlying ODE, and all results can be generalized to arbitrary probability flow ODE without loss of generality. We use the Fréchet inception distance (FID) to measure sampling quality on image data. Lower FID means better performance. See Appendix L for the list of full hyper-parameters employed in our numerical experiments. In addition,we provide some remarks on the training time consumption of our numerical experiments in Appendix M.
@@ 4.2.2. Real dataset
We compare the images generated by the numerical sampler and characteristic generator with different numbers of function evaluations (NFE) in Figure 4. The numerical sampler fails to accurately generate images at small NFE values such as 1 and 2. In fact, with 1 NFE, the solution is actually close to the mean of the target distribution. It is necessary to have five or more NFE for the numerical ODE solvers to work properly. In contrast, the characteristic generator is capable of generating high-quality images even with only 1 NFE. In essence, the characteristic generator distills the knowledge of a precise, multi-step solver into a fast, single step network. Our theory correctly predicts that its error should be much lower than that of a naive one-step Euler sampler, which aligns perfectly with our empirical findings in Figure 4. See Appendix O for detailed discusssion.
@@ 4.2.3. Comparison with other generative models
4.2.4. Self-Distillation on higher resolution
We also explore to apply the proposed method on higher resolution image datasets such as CelebA HQ (down-sampled to $256\times 256$ and $512\times 512$), and manage to eliminate the dependence on the teacher model, resulting a self-distillation scheme.
The main idea is to define a teacher-free reference solution $\overline{g}_{t,s}(X_{t})$ from a characteristic estimator $\widehat{g}$, by
(4.9) $\overline{g}_{t,s}(X_{t})=\widehat{g}_{u,s}\circ\widehat{g}_{t,u}(X_{t}),\ u\in(t,s).$
With the reference solution $\overline{g}$, we can get rid of the extra teacher network when training the characteristic generator, significantly reducing the required memory. Also, the formation of $\overline{g}$ naturally incorporates the semigroup constraint, as the optimization target is to minimize the distance between $\overline{g}$ and $\widehat{g}$. Notice that we do not specify the choice of $u$ in (4.9) as it should be arbitrary. Recent progress (Frans et al., 2025) finds $u=(t+s)/2$ works well enough and shows some symmetry, and we follow this setup. During the early stage of training, the local consistency loss plays the dominant role as the global counterpart relied on the reference solution, which requires the estimator $\widehat{g}$ fitted on a smaller time scale. During the late stage of training, the local consistency is already ensured, and the global consistency loss gradually takes effect.
We carry out the self-distillation variant on the CelebA HQ dataset (down-sampled to $256\times 256$ and $512\times 512$). We carry out the training and generation on the $8\times$ down-sampled latent space using the sd-vae-ft-mse autoencoder (Rombach et al., 2022). We also move from the U-Net architecture to the emerging DiT architectures (Peebles and Xie, 2023) (B/2 on the 256 resolution and XL/2 on the 512 resolution). We report the FID in Table 6 and the generated images in Figure 5.
Table 6. Comparison of FID by NFE by the characteristic generator on CelebAHQ-256 and CelebAHQ-512.
Resolution NFE=1 NFE=2 NFE=4 NFE=8
256 30.25 19.02 15.82 14.29
512 31.18 22.45 18.61 15.95
Figure 5. Generated images on CelebAHQ-256 (top) and CelebAHQ-512(bottom) (scaled down by a factor of 0.28 to fit the page).
@@ 6. Conclusions and Future Work
On the theoretical front, we intend to exploit the regularity of the velocity field in our error analysis for velocity matching. This will allow us to achieve a faster convergence rate. Additionally, we plan to analyze higher order and more stable numerical schemes, such as the high-order exponential integrator, in order to provide a comprehensive understanding of their effectiveness. Furthermore, we aim to establish a theoretical foundation for the role of semi-group penalties in the characteristic fitting. This will contribute to a deeper understanding of their impact and significance in our framework. A significant extension would be to move from our current average-in-time error guarantee (Theorem 3.14) to a stronger, uniform-in-time bound. Achieving this would likely require designing novel objective functionals for the characteristic fitting, potentially based on $L^{\infty}$-risk minimization or adversarial (minimax) formulations. Investigating the theoretical properties and empirical performance of these approaches is an important open question.
This work provides both theoretical analysis and experimental validation. However, it is important to clarify the role of our theoretical bounds, as their direct numerical verification is challenging due to the well-known gap between deep learning theory and practice. This gap arises from several fundamental factors: (i) Our bounds, like many in the field, serve to characterize scaling laws and qualitative relationships, e.g., how error scales with dimension, sample size, or the number of Euler steps. They are not intended to be numerically predictive of the exact empirical error. The bounds often contain large, abstract constants that are intractable to compute in practice, making a direct numerical comparison infeasible. (ii) Since the theoretical bounds often represent a worst-case scenario, there exists a gap between the empirical error and the theoretical bound. (iii) Our theory analyzes the properties of the empirical risk minimizer, whereas in practice, non-convex optimization methods find local minima. This creates a gap between the object of our theoretical analysis and the model obtained empirically. Bridging this theory-practice gap and developing bounds that more closely reflect the performance of practically trained models remains a crucial challenge for the field.
    Appendix A summarizes the notation used throughout the paper.
  2. (II)
    Appendix B provides auxiliary definitions and lemmas.
  3. (III)
    Appendix C presents the proofs for the theoretical results in Section 2.
  4. (IV)
    Appendix D establishes regularity properties of the probability flow ODE and provides the proofs for the propositions in Section 3.2.
  5. (V)
    Appendix E provides the proofs for the results in Section $3.$
  6. (VI)
    Appendix F provides the proofs for the results in Section 3.4.
  7. (VII)
    Appendix G provides the proofs for the results in Section 3.5.
  8. (VIII)
    Appendix H provides the proofs for the results in Section 3.7.
  9. (IX)
    Appendix I details a generalization error analysis for nonparametric regression, which is used to establish the oracle inequalities for velocity matching (Appendix E) and characteristic fitting (Appendix G).
  10. (X)
    Appendix J establishes approximation results for deep neural networks with a Lipschitz constraint, which are also used for the convergence rates of velocity matching (Appendix E) and characteristic fitting (Appendix G).
  11. (XI)
    Appendix K specifies the parameterization of the denoiser employed in Section 4.
  12. (XII)
    Appendix L reports the full hyperparameters used in our numerical experiments, including model size, training iterations, batch size, learning rate, optimizer, and augmentation choices.
  13. (XIII)
    Appendix M provides some remarks on the training time consumption of our numerical experiments.
  14. (XIV)
    Appendix N conducts a thorough comparison of the implementation details between our method and CTM.
  15. (XV)
    Appendix O conducts a thorough theoretical comparison between the one-step Euler sampler and characteristic generator.
Appendix A Summary of Notations
Table LABEL:tab:notations summarizes the notations used in throughout the paper for easy reference and cross checking.
Table 7. The list of notations used in Sections 2, and 3.
Symbols Description
$\mu_{0}$ Prior distribution, $\mu_{0}=N(0,I_{d})$ (Assumption 1).
$\mu_{1}$ Target distribution, $\mu_{1}=N(0,\sigma^{2}I_{d})*\nu$ (Assumption 2).
$\rho_{0},\rho_{1}$ Densities of $\mu_{0}$ and $\mu_{1}$.
$\mu_{t},\,\rho(t,x)$ Interpolant distribution and its density, see Prop. 2.1.
$(g_{t,s}^{*})_{\sharp}\mu_{t}$ Push-forward of $\mu_{t}$ by the flow $g_{t,s}^{*}$, exactly $\mu_{s}$2.1).
$\varphi_{d}$ Standard Gaussian density in ${\mathbb{R}}^{d}$ (Assumption 2).
$\nu$ Base distribution with $\operatorname{supp}(\nu)\subseteq[0,1]^{d}$ (Assumption 2).
$X_{0},X_{1}$ Independent samples from $\mu_{0}$ and $\mu_{1}$.
$X_{t}$ Stochastic interpolant $X_{t}=\alpha(t)X_{0}+\beta(t)X_{1}$ (Eq. (2.1)).
$\alpha(t),\beta(t)$ Interpolant coefficients (Cond. 1); write $\alpha_{t},\beta_{t}$.
$s^{*}(t,x)$ Score $\nabla\log\rho_{t}(x)$3.1).
$b^{*}(t,x)$ Velocity field (Eq. (2.3)); transport Eq. (2.2).
$g_{t,s}^{*}$ Continuous probability flow map: $x_{t}\mapsto x_{s}$2.1).
$Z_{t},Z_{s}$ Flowed variables $Z_{t}=g_{0,t}^{*}(Z_{0})$, $Z_{s}=g_{0,s}^{*}(Z_{0})$2.1).
$T$ Stopping time in $(1/2,1)$ used for training/evaluation.
${\mathcal{S}},\,n$ Training dataset for velocity matching $\{(t^{(i)},X_{0}^{(i)},X_{1}^{(i)})\}_{i=1}^{n}$ and its size.
${\mathcal{L}}(b)$ Population risk for velocity (Eq. (2.6)).
$\widehat{{\mathcal{L}}}_{n}(b)$ Empirical risk for velocity (Eq. (2.7)).
$\mathscr{B}$} Hypothesis class for velocity networks (Thm. 3.10).
$\widehat{b}$ Velocity estimator via ERM (Eq. (2.8)).
[…]
$\widehat{\mu}_{T}$ Push-forward at time $T$ under ODE with $\widehat{b}$3.3).
$K,\tau,t_{k}$ Number of steps, step size $\tau=T/K$, and grid $t_{k}=k\tau$2.3).
$\widehat{E}_{k,\ell}^{\tau}$ Euler flow map (Eq. (2.9)).
${\mathcal{Z}},\,m$ Teacher trajectories and their number for fitting (§2.4).
${\mathcal{R}}(g)$ Population characteristic fitting risk (Eq. (2.5)).
$\widehat{{\mathcal{R}}}_{T,m,K}^{\operatorname{Euler}}(g)$ Empirical characteristic risk using Euler data (Eq. (2.10)).
$\lambda,\,\widehat{{\mathscr{P}}}(g)$ Semi-group penalty weight and penalty (Eq. (2.12)).
$\widehat{{\mathcal{R}}}_{m,K}^{\operatorname{Euler},\lambda}(g)$ Penalized empirical risk (Eq. (2.12)).
$\Delta_{kj\ell}(g)$ Semi-group discrepancy term (§2.4).
$\mathscr{G}$} Hypothesis class for characteristic networks (Thm. 3.14).
$\widehat{g}_{t,s}$ Learned characteristic generator (ERM; Eq. (2.11)).
[…]
$B_{\operatorname{vel}},B_{\operatorname{flow}},B^{\prime}_{\operatorname{flow}}$ Local bounds for velocity/flow and time-derivatives (§3.2).
$G$ Uniform bound on spatial Jacobians (Prop. 3.5).
$D$ Time-regularity constant for $\partial_{t}b^{*}$ (Prop. 3.8).
$\kappa(T)$ Function in time-regularity bounds (Prop. 3.8).
$C$ Universal constants depending on $d,\sigma$ (various theorems).
$d^{*}$ Intrinsic dimension with $d^{*}\ll d$3.7).
$P\in{\mathbb{R}}^{d\times d^{*}}$ Column-orthonormal embedding matrix (§3.7).
$\widetilde{\nu},\,P_{\sharp}\widetilde{\nu}$ Base distribution on $0,1]^{d^{*}}$ and its push-forward to ${\mathbb{R}}^{d}$.
$\widetilde{b}^{*},\,\widetilde{g}^{*}$ Intrinsic-coordinate velocity/flow (Lemma 3.18).
Appendix B Supplemental Definitions and Lemmas
@@
We first introduce the sub-Gaussian random variable Vershynin [2018], Wainwright [2019].
Definition B.1 (Sub-Gaussian).
@@
The following results (Lemmas B.2 to B.4) shows some important properties of sub-Gaussian variables, whose proofs can be found in [Wainwright, 2019, Section 2.1].
Lemma B.2 (Chernoff bound).
@@
Lemmas B.5 and B.6 show bounds of the tail probability and the expectation of the maximum of $N$ sub-Gaussian variables.
Lemma B.5.
@@
where the last inequality holds from Lemma B.2. This completes the proof. ∎
Proof of Lemma B.6.
@@
where the last inequality holds from Lemma B.3 for each $\lambda\in[0,1)$. Letting $\lambda=1/2$ yields the desired inequality. ∎
Lemma B.7 (Fourth moment of standard Gaussian).
@@
We next introduce the notation of covering number and Vapnik-Chervonenkis dimension (VC-dimension), both of which measure the complexity of a function class Mohri et al. [2018], Vaart and Wellner [2023]. They are used in the error analysis for velocity matching (Section E) and characteristic fitting (Section G).
Definition B.8 (Covering number).
@@
Definition B.9 (VC-dimension).
Let $\mathscr{F}$} be a class of functions from ${\mathbb{R}}^{d}$ to $\{\pm 1\}$. For any non-negative integer $m$, we define the growth function of $\mathscr{F}$} as
@@
A set $\{X^{(i)}\}_{i=1}^{m}$ is said to be shattered by $\mathscr{F}$} when $|\{(f(X^{(1)}),\ldots,f(X^{(m)})):f\in{\mathscr{F}}\}|=2^{m}$. The Vapnik-Chervonenkis dimension of $\mathscr{F}$}, denoted $\operatorname{VCdim}({\mathscr{F}})$, is the size of the largest set that can be shattered by $\mathscr{F}$}, that is, $\operatorname{VCdim}({\mathscr{F}})=\max\{m:\Pi_{{\mathscr{F}}}(m)=2^{m}\}$. For a class $\mathscr{F}$} of real-valued functions, we define $\operatorname{VCdim}({\mathscr{F}})=\operatorname{VCdim}(\operatorname{sign}({\mathscr{F}}))$.
Lemma B.10.
Let $\mathscr{F}$} be a class of functions mapping from ${\mathbb{R}}^{d}$ to $\mathbb{R}$}, and let $\mathscr{H}$} be a function class defined as ${\mathscr{H}}=\{(x,f)\mapsto h(f,x)\in{\mathbb{R}}:f\in{\mathscr{F}}\}$. Suppose ${\mathcal{D}}=\{X^{(i)}\}_{i=1}^{n}$ is a set of samples in ${\mathbb{R}}^{d}$. If there exists a constant $L>0$ such that for each $f,f^{\prime}\in{\mathscr{F}}$,
@@
Lemma B.11 ([Anthony et al., 1999, Theorem 12.2]).
@@
Lemma B.12 ([Bartlett et al., 2019, Theorem 7]).
@@
Appendix C Proof of Results in Section 2
The proof of Proposition 2.1 follows from the proof of [Albergo et al., 2023b, Theorem 2] and [Albergo et al., 2023a, Theorem 2.6].
@@ Appendix C Proof of Results in Section 2
(C.1) $\frac{\partial}{\partial t}\Big(\int_{{\mathbb{R}}^{d}}\phi(x)\rho_{t}(x)\,\mathrm{d}x\Big)=\int_{{\mathbb{R}}^{d}}\phi(x)\partial_{t}\rho_{t}(x)\,\mathrm{d}x.$
@@
(C.2) $=\int_{{\mathbb{R}}^{d}}b_{t}^{*}(x)\cdot\nabla\phi(x)\rho_{t}(x)\,\mathrm{d}x=-\int_{{\mathbb{R}}^{d}}\phi(x)\nabla\cdot(b_{t}^{*}(x)\rho_{t}(x))\,\mathrm{d}x,$
where the first equality follows from the chain rule, the forth equality holds from the definition of $b_{t}^{*}$ (2.3), and the last equality is due to integration by parts and the divergence theorem [Evans, 2010, Theorem 1 in Section C.2]. Combining (C.1) and (C.2) gives the desired transport equation. ∎
Appendix D Properties of the Probability Flow ODE
In this section, we present some auxiliary properties of the probability flow ODE in Section D.1. Proofs of the propositions in Section 3.2 are given in Section D.2.
D.1. Auxiliary Properties
Lemmas D.1 to D.4 are established under Assumption 1.
Lemma D.1.
@@ D.1. Auxiliary Properties
where the we used the definition of conditional density and Lemma D.1. This completes the proof. ∎
Lemma D.3.
@@
(D.1) $b^{*}(t,x)=\frac{\dot{\alpha}_{t}}{\alpha_{t}}x+\beta_{t}\Big(\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\Big){\mathbb{E}}[X_{1}|X_{t}=x],$
@@
(D.2) $\nabla b^{*}(t,x)=\frac{\dot{\alpha}_{t}}{\alpha_{t}}I_{d}+\beta_{t}\Big(\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\Big)\nabla{\mathbb{E}}[X_{1}|X_{t}=x].$
@@
(D.3) $=-\nabla_{x}\log\rho_{t}(x)\int_{{\mathbb{R}}^{d}}x_{1}^{T}\rho_{1|t}(x_{1}|x)\,\mathrm{d}x_{1}+\int_{{\mathbb{R}}^{d}}\nabla_{x}\log\rho_{t|1}(x|x_{1})x_{1}^{T}\rho_{1|t}(x_{1}|x)\,\mathrm{d}x_{1},$
@@
For the first term in the right-hand side of (D.3), applying Lemma D.2 implies
@@
(D.4) $=\Big(\frac{1}{\alpha_{t}^{2}}x-\frac{\beta_{t}}{\alpha_{t}^{2}}{\mathbb{E}}[X_{1}|X_{t}=x]\Big){\mathbb{E}}[X_{1}|X_{t}=x]^{T}.$
For the second term, it follows from Lemma D.1 that
@@
(D.5) $=-\frac{1}{\alpha_{t}^{2}}x{\mathbb{E}}[X_{1}|X_{t}=x]^{T}+\frac{\beta_{t}}{\alpha_{t}^{2}}{\mathbb{E}}[X_{1}X_{1}^{T}|X_{t}=x].$
Plugging (D.4) and (D.5) into (D.3) yields
(D.6) $\nabla{\mathbb{E}}[X_{1}|X_{t}=x]=\frac{\beta_{t}}{\alpha_{t}^{2}}\operatorname{Cov}(X_{1}|X_{t}=x).$
Substituting (D.6) into (D.2) completes the proof. ∎
@@
Lemmas D.5 to D.8 are established under Assumptions 1 and 2.
Lemma D.5.
@@
$(X_{1}|X_{t}=x)\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}U_{t,x}+\sqrt{\frac{\sigma^{2}\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}}\epsilon+\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}x,$
where $U_{t,x}$ is a random variable satisfying $\operatorname{supp}(U_{t,x})=\operatorname{supp}(\nu)$, and $\epsilon\sim N(0,I_{d})$.
Proof of Lemma D.5.
@@
(D.7) $\rho_{t|1}(x|x_{1})=\frac{1}{(2\pi)^{d/2}\alpha_{t}^{d}}\exp\Big(-\frac{1}{\alpha_{t}^{2}}\|\beta_{t}x_{1}-x\|_{2}^{2}\Big).$
@@
(D.8) $\rho_{1}(x_{1})=\frac{1}{(2\pi)^{d/2}\sigma^{d}}\int_{{\mathbb{R}}^{d}}\exp\Big(-\frac{1}{\sigma^{2}}\|x_{1}-u\|_{2}^{2}\Big)\,\mathrm{d}\nu(u).$
Combining (D.7) and (D.8) implies